Operations Report – March 8th

Crew 223 Operations Report 08/03/2020

SOL: 7

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report
Notes on non-nominal systems: Nothing to report

Saturday generator:
Off at 6:00 am
On at 8:57 pm SOC was 86%

Sunday generator:
Off at 7:30 am
On at 9:00 pm SOC was XX%

Diesel Reading – 68%
Station Propane Reading – 75%
Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): 47 gallons
Water Meter: 149748.04 units
Water (static tank) – 280 gallons
Static to Loft Pump used – yes
Water in GreenHab – 135.2 gallons
Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Perseverance rover used: yes
Hours: 150.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 134.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.5
Beginning charge: 100%
Ending charge: 96%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)

Reason for use: I started the 4 ATVs during the engineer EVA for battery check. I had trouble starting the Honda ATV (same as sol 4). It sounds like a flat battery. We will try again to start it during the EVA tomorrow, and if not starting we will use the jumper cables or bring back the battery to charge it in the RAM.

Oil Added? no

ATV Fuel Used: Nothing to report

Hours the ATVs were Used today: Nothing to report.

Notes on ATVs: Nothing to report.

HabCar used and why, where? The HabCar is parked at the MDRS Campus.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.
Summary of suits and radios: Suits 7, 10 were used during the EVA.
Summary of Hab operations: Nothing to report.
Summary of GreenHab operations: Nothing to report.
Summary of ScienceDome operations: Nothing to report.
Summary of RAM operations: Nothing to report.
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing to report

Operations Report – March 7th

Crew 223 Operations Report 07/03/2020

SOL: 6

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Friday generator:
Off at 6:00am
On at 8:53pm SOC was 88%

Saturday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 75%

Ethanol Free Gasoline – 5.5 gallons (1.5 taken today to fill up the ATVs)

Water (loft tank): 47 gallons

Water Meter: 149736.10 units

Water (static tank) – 292 gallons

Static to Loft Pump used – yes

Water in GreenHab – 143.1 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: yes
Hours: 150.1
Beginning charge: 100%
Ending charge: 97%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes
Hours: 134.1
Beginning charge: 100%
Ending charge: 63%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 139.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: I started the 4 ATVs during the engineer EVA for battery
check.
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: Nothing to report.
Notes on ATVs: Thanks to Shannon, all the ATVs are full of gas.

HabCar used and why, where? The HabCar is parked at the MDRS Campus.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 7, 8, 9 were used during the EVA.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 06th

  

Crew 223 Operations Report 06/03/2020

SOL: 5

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Thursday generator:
Off at 4:30am
On at 9:10pm SOC was 84%

Friday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 75%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149721.59 units

Water (static tank) – 307 gallons

Static to Loft Pump used – yes

Water in GreenHab – 153.7 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: yes
Hours: 149.7
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 133.7
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.5
Beginning charge: 100%
Ending charge: 80%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Honda started during the EVA this morning, and we had no
problem with it. I started the 4 ATVs during the engineer EVA for
battery check.
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? The HabCar is parked at the MDRS Campus.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.
No problem.

Summary of Hab operations: One of the gaz button of the gaz stove is
broken (the one for the upper left).

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Florian built the solar hoven.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 05th

Crew 223 Operations Report 05/03/2020

SOL: 4

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Wednesday generator:
Off at 6:00am
On at 9:15pm SOC was 84%

Thursday generator:
Off at 4:30am
On at 9pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149709.09 units

Water (static tank) – 319 gallons

Static to Loft Pump used – yes

Water in GreenHab – 166.4 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Perseverance (former Demos) rover used: no
Hours: 149.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes
Hours: 133.7
Beginning charge: 100%
Ending charge: 86%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 139.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: 1 & 2 used during the EVA. We had trouble starting the
Honda, and finally thought the battery was flat. But during the engineer
EVA, Honda started without any problem (using the starter). So maybe it
was due to the cold weather… Crew members going in EVA tomorrow
morning will try to start it. If non working, we will probably take the
battery in the RAM to charge it. All ATVs were run during the engineer
EVA for battery check.
Oil Added? no
ATV Fuel Used: ~5%
Hours the ATVs were Used today: 20 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? Nothing to report

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 8, 9, 10 were used during the EVA.
Suit 7 is now fixed and perflectly working !

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: The GreenHab officier switched the
working and non working neons of the halogen lamp in order to have more
light on tomatoes than on aromatics.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: We successfully fixed suit 7 ! What we found
is a problem on the suit female plug : in a certain position, one of the
cables was in contact with one of the 2 free pins. And thus connecting
the electrical circuit to the ground. Our solution was to isolate those
2 pins with an heat-shrinkable sleeve. Cf the 2 pictures attached. We
did the same thing on suits 8, 9, 10 to prevent them from this.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 4th

Crew 223 Operations Report 04/03/2020

SOL: 3

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Tuesday generator:
Off at 5:00am
On at 9:10pm SOC was 84%

Wednesday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149694.75 units

Water (static tank) – 334 gallons

Static to Loft Pump used – yes

Water in GreenHab – 180.2 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 149.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover used: no
Hours: 133.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.3
Beginning charge: 100%
Ending charge: 91%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? Nothing to report

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 2, 3, 4, 9 were used during the EVA.
Suit 10 is working, and we will use it tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: We continue to fix suit 7 but haven’t
finished now. We first found a loose electrical connection at the
connection with the battery at the minus pod. We corrected it. But there
is still a problem with the suit, and we tend for another loose
electrical connection in the control box.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 3rd

Crew 223 Operations Report 03/03/2020

SOL: 3

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Monday generator:
Off at 4:30am
On at 9:10pm SOC was 83%

Tuesday generator:
Off at 5:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149678.53 units

Water (static tank) – 350 gallons

Static to Loft Pump used – yes

Water in GreenHab – 201.4 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 149.4
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes
Hours: 133.6
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 139.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: The ATV 1 and 2’s oil light up when starting the ATV and
then turn off.

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 8, 9, 10, 2 were used during the
EVA. We had troubles with suit number 10 : the fans stopped after 26
minutes. We think this is because the suit was not well plugged and thus
was not charged. We are checking this and will continue tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Following Scott Davis’ recommendations, we
tried to fix suit 7. We checked most of the electrical connection points
and it doesn’t seem to be a loose electrical connection. We tend for a
problem with the suit’s battery (but we need to confirm this guess
tomorrow), and we will also remove the fans to make sure they are not
obstructed.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Do you know were is
the key of the upper black storage box in the RAM (I would like to open
the drawer containing the wire strippers) ? I only found the one opening
the lower black storage box.

Operations Report – March 1st

Crew 223 Operations Report 03/01/2020

SOL: 0

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Saturday generator:
Off at 5:30am
On at 8:00pm SOC was 80%

Sunday generator:
Off at 5:30am
On at 7:00pm SOC was 79%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149647.32 units

Water (static tank) – 381 gallons

Static to Loft Pump used – yes

Water in GreenHab – 235.2 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (for training with Shannon)
Hours: 147
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes (for training with Shannon)
Hours: 133.2
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes (for training with Shannon)
Hours: 139.1
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Training with Shannon
Oil Added? no
ATV Fuel Used: almost 2% of the tank.
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: As it already happened with the previous crew, the Honda ATV was stuck in 5th gear and refused to go back into neutral (reason = slowing down while 5th gear engaged). Shannon fixed it by moving manually the ATV. The gears are now working perfectly.

HabCar used and why, where? We used the Habcar to go to Hanksville to fill up the GreenHab tank (which was full at noon). We also refueled it. Habcar gas tank cap broken. The HabCar is now parked at the MDRS.

CrewCar used and why, where? The CrewCar is in Grand Junction since this morning, thanks to crew 222.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suit 3 was used during the engineer EVA.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 2nd

Crew 223 Operations Report 03/02/2020

SOL: 1

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Sunday generator:
Off at 5:30am
On at 7:00pm SOC was 79%

Monday generator:
Off at 4:30am
On at 9pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149665.40 units

Water (static tank) – 363 gallons

Static to Loft Pump used – yes

Water in GreenHab – 218.3 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: yes
Hours: 149.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 133.2
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: The ATV 1’s oil light up.

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 7, 8, 9 were used during the EVA. We had troubles with suit number 7 (intermittent fan, same as crew 222 and 221), I tried to fix it but didn’t manage to. I just find an email sent last friday by Scott Davis giving some ideas to fix it, so I will try again tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Tried to fix suit 7 without success.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 2nd

Crew 223 Operations Report 03/01/2020

SOL: 0

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Saturday generator:
Off at 5:30am
On at 8:00pm SOC was 80%

Sunday generator:
Off at 5:30am
On at 7:00pm SOC was 79%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149647.32 units

Water (static tank) – 381 gallons

Static to Loft Pump used – yes

Water in GreenHab – 235.2 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (for training with Shannon)
Hours: 147
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes (for training with Shannon)
Hours: 133.2
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes (for training with Shannon)
Hours: 139.1
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Training with Shannon
Oil Added? no
ATV Fuel Used: almost 2% of the tank.
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: As it already happened with the previous crew, the Honda ATV was stuck in 5th gear and refused to go back into neutral (reason = slowing down while 5th gear engaged). Shannon fixed it by moving manually the ATV. The gears are now working perfectly.

HabCar used and why, where? We used the Habcar to go to Hanksville to fill up the GreenHab tank (which was full at noon). We also refueled it. Habcar gas tank cap broken. The HabCar is now parked at the MDRS.

CrewCar used and why, where? The CrewCar is in Grand Junction since this morning, thanks to crew 222.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suit 3 was used during the engineer EVA.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Operations Report – March 1st

Crew 223 Operations Report 03/01/2020

SOL: 0

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Saturday generator:
Off at 5:30 am
On at 8:00 pm SOC was 80%

Sunday generator:
Off at 5:30 am
On at 7:00 pm SOC was 79%

Diesel Reading – 70%
Station Propane Reading – 82%
Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons
Water Meter: 149647.32 units
Water (static tank) – 381 gallons
Static to Loft Pump used – yes
Water in GreenHab – 14.8 gallons
Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (for training with Shannon)
Hours: 147
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes (for training with Shannon)
Hours: 133.2
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes (for training with Shannon)
Hours: 139.1
Beginning charge: 100%
Ending charge: Non checked
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Training with Shannon
Oil Added? no
ATV Fuel Used: almost 2% of the tank.
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: As it already happened with the previous crew, the Honda
ATV was stuck in 5th gear and refused to go back into neutral (reason = slowing down while 5th gear engaged). Shannon fixed it by moving
manually the ATV. The gears are now working perfectly.

HabCar used and why, where? We used the Habcar to go to Hanksville to fill up the GreenHab tank (which was full at noon). We also refueled it.
Habcar gas tank cap broken. The HabCar is now parked at the MDRS.

CrewCar used and why, where? The CrewCar is in Grand Junction since this morning, thanks to crew 222.

General notes and comments: Nothing to report

Summary of the internet: All ok.

Summary of suits and radios: Suit 3 was used during the engineer EVA.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing to report